3 You are considering starting a greenhouse tomato production facility. You estimate your total variable costs...

3

  1. You are considering starting a greenhouse tomato production facility. You estimate your total variable costs will be $20,000 per greenhouse and your total fixed costs will be $15,000 per greenhouse. You anticipate receiving a price of $1.25/pound for your tomatoes. What is the minimum number of pounds of tomatoes you need to sell to be profitable in the long-run? If you were already producing, what is the minimum number of pounds of tomatoes you need to sell to continue producing in the short-run?
Answer #1

Total cost will be=20000+15000=35000.

To do profit minimum TR should be more than 35000.

So ww beed to divide the TR by price to get quantity

=35000/1.25

=28000 unit

So here minimum 28000 should be produced to earn zero economic profit.

If you were already producing, what is the minimum number of pounds of tomatoes you need to sell to continue producing in the short-run?

Answer:

Firm shutdowns when Price is below AVC.

AVC is average variable cost.

So, firm has to produce maximum tomato that covers minimum cost.

To find that quantity we must divide the total variable cost by price.

=20000/1.25

=16000.

So minimum 16000 quantity should be produce to run the business.